OpenStudy (anonymous):

y(x) = ln[sqrt(2 − x^2)] what is y'(x)?

OpenStudy (anonymous):

\[f(x)=\ln(g(x))\] \[f'(x)=\frac{g'(x)}{g(x)}\]

OpenStudy (anonymous):

thanks, that's useful to know!! is that rule called anything in particular?

OpenStudy (anonymous):

oh i have a better idea, i wasn't paying attention

OpenStudy (anonymous):

\[f(x)=\ln(\sqrt{2-x^2})\] \[f(x)=\frac{1}{2}\ln(2-x^2)\]take the derivative now

OpenStudy (anonymous):

it is the chain rule for logs and it is very useful

OpenStudy (anonymous):

it is always the case that \[\frac{d}{dx}\ln(g(x))=\frac{g'(x)}{g(x)}\]

OpenStudy (anonymous):

Can you follow the instruction?